OpenStudy (anonymous):

can u help me @matthew1286 or @clynnew

OpenStudy (anonymous):

can u help me please

OpenStudy (anonymous):

Oliver’s room is 3 yards wide. The perimeter of the room is 14 yards. What is the area of Oliver’s room?

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

do you know how to find the area?

OpenStudy (anonymous):

multiple the width with the hight

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

yes im here. so 3 * 3 = ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

okay and do you have any choics for this problem?

OpenStudy (anonymous):

no so whats the answer

OpenStudy (anonymous):

well try this problem \[\frac{ 1 }{ 16} P ^{2}\]

OpenStudy (anonymous):

is 9 the answer

OpenStudy (anonymous):

and instead of P plug in your perimeter.. of 14.

OpenStudy (anonymous):

what goes beyween the p x or +

OpenStudy (anonymous):

\[( \frac{ 1 }{ 16 } ) (\frac{ 14 }{ 1 })^{2}\]

OpenStudy (anonymous):

solve that.

OpenStudy (anonymous):

49/4

OpenStudy (anonymous):

can you show me your work?

OpenStudy (anonymous):

k let me get it

OpenStudy (anonymous):

on how i solve it

OpenStudy (anonymous):